OpenStudy (berrymox):

Show the following limit:

OpenStudy (berrymox):

\[\lim_{h \rightarrow 0}\frac{ 1 }{ \sqrt{h+1} +1} = 0.5\]

OpenStudy (agent0smith):

Multiply the numerator and denominator by the conjugate of the denominator, which is \[\large\sqrt{h+1} -1\]

OpenStudy (berrymox):

The furthest I got was 1 /[ (sqrt(h+1) -1 ]

OpenStudy (berrymox):

I don't know where to go from here

OpenStudy (berrymox):

oh wait

OpenStudy (berrymox):

I always don't know when to plug-in because I assume I get 0/0 or i >_<

OpenStudy (berrymox):

thank you

OpenStudy (michele_laino):

it is simple, the function: \[f\left( h \right) = \frac{1}{{\sqrt {h + 1} + 1}}\] is continuous at \(h=0\), so please substitute \(h=0\), what do you get?

OpenStudy (agent0smith):

Oh yeah it's not even -1. Just plug it in as Michele said.

OpenStudy (agent0smith):

What I gave you is for if the function was 1 /[ (sqrt(h+1) -1 ]

OpenStudy (agent0smith):

"I always don't know when to plug-in because I assume I get 0/0 or i" If you don't always know... then plug in and find out. In this case if you plug in h=0, you'll discover it is not 0/0.

OpenStudy (berrymox):

It's because I try to over simplify, :P